# Can Test-Time Scaling Improve World Foundation Model?

### v1

@@ Abstract
World foundation models, which simulate the physical world by predicting future states from current observations and inputs, have become central to many applications in physical intelligence, including autonomous driving and robotics. However, these models require substantial computational resources for pretraining and are further constrained by available data during post-training. As such, scaling computation at test time emerges as both a critical and practical alternative to traditional model enlargement or re-training. In this work, we introduce SWIFT, a test-time scaling framework tailored for WFMs. SWIFT integrates our extensible WFM evaluation toolkit with process-level inference strategies, including fast tokenization, probability-based Top-K pruning, and efficient beam search. Empirical results on the COSMOS model demonstrate that test-time scaling exists even in a compute-optimal way. Our findings reveal that test-time scaling laws hold for WFMs and that SWIFT provides a scalable and effective pathway for improving WFM inference without retraining or increasing model size. The code is available at https://github.com/Mia-Cong/SWIFT.git.
@@ 1 Introduction
Figure 1: Can Test-Time Scaling Improve World Foundation Model? We propose the first test-time scaling framework tailored for WFMs, SWIFT, and provide a testbed for evaluation.
@@ 3 WFM Evaluation Toolkit
Figure 2: Videos of world foundation model without (top) and with (bottom) our test-time scaling framwork.
@@ 4.3 Scaling via Test-Time Verifiers: Action $\mathcal{A}$ Design
Figure 4: ORM, PRM, and Beam Search.
@@ 4.4.1 Naive best-of-N to verify test-time scaling law
We begin with the simplest possible test‑time scaling strategy: a best‑of‑$N$ search. For each input, we sample $N$ independent video continuations, fully generate each one, score them using our rule‑based verifiers $\mathcal{R}$, and then select the highest‑scoring video (Figure 4, left). This approach mirrors the “best‑of‑$N$” output selection commonly used in LLM research (ORM). Although conceptually trivial, best‑of‑$N$ serves as a crucial start to investigate test-time scaling laws in WFM.
@@ 4.4.2 Adapting Test-Time Scaling Strategy to World Modeling Practically and Efficiently
Challenge: Expensive decoding via a diffusion-based decoder: WFMs rely on a diffusion‑based denoising tokenizer to produce final high‑quality frames — a process that takes $\approx$ 137.2 s on A6000, roughly as long as autoregressively generating those frames. Hence, performing a full diffusion decode
Figure 5: Scores on videos decoded by diffusion decoder (slow) and tokenizer decoder (fast) follow the same trend.
step-wise creates substantial overhead. While test‑time scaling inherently trades extra compute for better output quality, an efficient strategy must avoid wasted decoding and instead focus compute on exploring promising candidate trajectories.
@@
1. Fast-Tokenizer to Accelerate Decision Process  To avoid running the diffusion decoder at every generation step, we leverage the WFM’s discrete-token decoder as a lightweight proxy. This “fast tokenizer” converts intermediate latent outputs into video frames in only $\approx$0.015 s — versus $\approx$130 s for a full diffusion decode. We find that the reward scores computed from fast-tokenized outputs show strong correlation with those from the diffusion decoder (see Figure 5). This tight alignment confirms that the fast tokenizer offers a reliable and drastically cheaper feedback signal, enabling early pruning of low-quality trajectories and allowing compute to focus on more promising candidate paths.
@@ 5.1 Quantitative Results
Table 1: Comparison of COSMOS-4B and COSMOS-12B using naive best-of-$N$ sampling. FVD and FID scores are reported separately for the nuScenes (left) and Waymo (right) datasets, due to significant differences in their underlying data distributions.
@@
We also investigate the text+image-to-video modality; see Appendix C for more details.
Model N Alg. FVD FID IQ MS VQ TC DD FC
@@
Table 2: Comparison between COSMOS-4B and COSMOS-12B under different search algorithms. $N$ is sample number. For our beam search with probability, $M$ is set as sqrt($N$).
@@
Human Evaluation. To better demonstrate the performance gain over larger pretrained model, we opt for human evaluation for more faithful analysis. Following recent advances Bar-Tal et al. (2024); Blattmann et al. (2023a); Chen et al. (2023), we adopt the Two-Alternative Forced Choice (2AFC) protocol. Specifically, participants are shown side-by-side video pairs and asked to select the one they find better. We collect a total of 1,885 responses from 12 participants. As shown in Figure 8, outputs from the smaller model enhanced with test-time scaling are often preferred over those from the larger baseline. While this study focuses on scaling smaller models, we note that applying test-time scaling to larger models can further push performance toward the frontier—an exciting direction we leave for future work.
@@ 5.2 Qualitative Results
We present qualitative examples in Figures 9 and 10 (Appendix B). The baseline model frequently produces inconsistent objects, unstable backgrounds, or unnatural appearances. In contrast, test-time scaling consistently mitigates these issues—enhancing temporal consistency and improving overall visual fidelity.
@@ Appendix A More Related Works
Appendix B Qualitative Results
We present visualizations of videos generated by COSMOS-4B and COSMOS-5B in Figures 9 and 10, respectively. Each example includes video outputs under different test-time scaling strategies. The results illustrate that both video quality and text alignment consistently improve with test-time scaling, especially when using our efficient and reliable inference strategies.
Appendix C Text+image to Video Modality Evaluation.
We also prove our findings with the text+image to video modality and report the results in Table 3. The findings are aligned with those on the image-to-video modality.
Model N Alg. FVD FID IQ MS VQ TC DD FC
5B 1 - 728.68/126.63 59.71/10.47 63.48 0.9822 3.80 3.44 3.81 3.63
2 ORM 659.79/111.96 59.45/10.01 63.90 0.9840 3.89 3.55 3.88 3.71
4 PRM 641.5/112.15 52.60/10.07 64.39 0.9843 3.89 3.55 3.88 3.71
Ours 628.01/110.02 50.16/9.98 64.77 0.9848 3.91 3.57 3.92 3.73
13B 1 - 616.92/109.77 59.71/11.48 63.79 0.9834 3.92 3.55 3.94 3.75
Table 3: Comparison between COSMOS-5B and COSMOS-13B under different search algorithms. $N$ is sample number. For our beam search with probability, $M$ is set as sqrt($N$).

### v2

@@ Abstract
World foundation models, which simulate the physical world by predicting future states from current observations and inputs, have become central to many applications in physical intelligence, including autonomous driving and robotics. However, these models require substantial computational resources for pretraining and are further constrained by available data during post-training. As such, scaling computation at test time emerges as both a critical and practical alternative to traditional model enlargement or re-training. In this work, we introduce SWIFT, a test-time scaling framework tailored for WFMs. SWIFT integrates our extensible WFM evaluation toolkit with process-level inference strategies, including fast tokenization, probability-based Top-K pruning, and efficient beam search. Empirical results on the COSMOS model demonstrate that test-time scaling exists even in a compute-optimal way. Our findings reveal that test-time scaling laws hold for WFMs and that SWIFT provides a scalable and effective pathway for improving WFM inference without retraining or increasing model size. Project page: https://scalingwfm.github.io/.
@@ 1 Introduction
Figure 1: The SWIFT pipeline for studying test-time scaling (TTS) in world foundation models (WFMs). We use autoregressive world model COSMOS as a base, which initially produces an unrealistic world simulation (top panel). By applying our TTS method, the simulation is significantly enhanced and becomes more physically plausible (bottom panel).
@@ 3 WFM Evaluation Toolkit
Figure 2: Videos of world foundation model without (top) and with (bottom) TTS, where TTS improves 3D consistency (left) and temporal consistency (right) in the generated videos.
@@ 4.3 Scaling via Test-Time Verifiers: Action $\mathcal{A}$ Design
Figure 4: Compared to ORM and PRM, our proposed beam search with probability is more efficient and practical for WFMs with three key efficiency designs.
@@ 4.4.1 Naive best-of-N to verify test-time scaling law
We begin with the simplest possible test‑time scaling strategy: a best‑of‑$N$ search. For each input, we sample $N$ independent video continuations, fully generate each one, score them using our rule‑based verifiers $\mathcal{R}$, and then select the highest‑scoring video (Figure 4, left). This approach mirrors the “best‑of‑NN” output selection commonly used in LLM research (ORM). Although conceptually trivial, best‑of‑$N$ serves as a crucial start to investigate test-time scaling laws in WFM.
@@ 4.4.2 Adapting Test-Time Scaling Strategy to World Modeling Practically and Efficiently
Challenge: Expensive decoding via a diffusion-based decoder: WFMs rely on a diffusion‑based denoising tokenizer to produce high‑quality frames. This process takes $\approx$ 137.2 s on A6000, roughly as long as autoregressively generating those frames. Hence, performing a full diffusion decode step-wise creates substantial overhead. While test‑time scaling inherently trades extra compute for better output quality, an efficient strategy must avoid wasted decoding and instead focus compute on exploring promising candidate trajectories.
@@
1. Fast-Tokenizer to Accelerate Decision Process To avoid running the diffusion decoder at every generation step, we leverage the WFM’s discrete-token decoder
Figure 5: Scores for videos decoded by diffusion decoder (slow) and tokenizer decoder (fast) follow the same trend.
as a lightweight proxy. This “fast tokenizer” converts intermediate latent outputs into video frames in only $\approx$0.015 s — versus $\approx$130 s for a full diffusion decode. We find that the reward scores computed from fast-tokenized outputs show strong correlation with those from the diffusion decoder (see Figure 5). This tight alignment confirms that the fast tokenizer offers a reliable and drastically cheaper feedback signal, enabling early pruning of low-quality trajectories and allowing compute to focus on more promising candidate paths.
@@ 5.1 Quantitative Results
Table 1: Comparison of COSMOS-4B and COSMOS-12B using naive best-of-NN sampling. FVD and FID scores are reported separately for the nuScenes (left) and Waymo (right) datasets, due to significant differences in their underlying data distributions.
@@
Model N Alg. FVD FID IQ MS VQ TC $D$ FC
@@
Table 2: Comparison between COSMOS-4B and COSMOS-12B under different search algorithms. NN is sample number. For our beam search with probability, $M$ is set as sqrt(NN).
We also investigate the text+image-to-video modality in Table 3 and the findings are aligned with those on the image-to-video modality.
Model N Alg. FVD FID IQ MS VQ TC $D$ FC
5B 1 - 728.68/126.63 59.71/10.47 63.48 0.9822 3.80 3.44 3.81 3.63
2 ORM 659.79/111.96 59.45/10.01 63.90 0.9840 3.89 3.55 3.88 3.71
4 PRM 641.5/112.15 52.60/10.07 64.39 0.9843 3.89 3.55 3.88 3.71
Ours 628.01/110.02 50.16/9.98 64.77 0.9848 3.91 3.57 3.92 3.73
13B 1 - 616.92/109.77 59.71/11.48 63.79 0.9834 3.92 3.55 3.94 3.75
Table 3: Comparison between COSMOS-5B and COSMOS-13B under different search algorithms. NN is sample number. For our beam search with probability, $M$ is set as sqrt(NN).
@@
Human Evaluation. To better demonstrate the performance gain over larger pretrained model, we opt for human evaluation for more faithful analysis. Following recent advances Bar-Tal et al. (2024); Blattmann et al. (2023a); Chen et al. (2023), we adopt the Two-Alternative Forced Choice (2AFC) protocol. Specifically, participants are shown side-by-side video pairs and asked to select the one they find better. We collect a total of 3,685 responses from 24 participants. As shown in Figure 8, outputs from the smaller model enhanced with test-time scaling are often preferred over those from the larger baseline. While this study focuses on scaling smaller models, we note that applying test-time scaling to larger models can further push performance toward the frontier—an exciting direction we leave for future work.
@@ 5.2 Qualitative Results
We present qualitative examples in Figure 9 and Figure 10. Each example includes video outputs under different test-time scaling strategies. The baseline model frequently produces inconsistent objects, unstable backgrounds, or unnatural appearances. In contrast, test-time scaling consistently mitigates these issues—enhancing temporal consistency and improving overall visual fidelity.